### v3

# A Constant Factor Approximation for Directed Feedback Vertex Set in Graphs of Bounded Genus

Abstract
The minimum directed feedback vertex set problem consists in finding the minimum set of vertices that should be removed in order to make a directed graph acyclic. This is a well-known NP-hard optimization problem with applications in various fields, such as VLSI chip design, bioinformatics and transaction processing deadlock prevention and node-weighted network design. We show a constant factor approximation for the directed feedback vertex set problem in graphs of bounded genus.
1 Introduction
In the directed feedback vertex set problem (DFVS), we are given a directed graph $G=(V,E)$ and costs $c_{v}\ \forall v\in V$ and wish to find a minimum cost set $X$ for which $G\backslash X$ is acyclic. DFVS is one of Karp’s original 21 NP-hard problems [15]. The DFVS problem has many applications including deadlock resolution [10], VLSI chip design [18] and program verification [19].
A 2-approximation for (undirected) FVS is given in [2]. DFVS has a 2-approximation in tournaments [20] and bipartite tournaments [23], is polynomial-time solvable on graphs of bounded treewidth, has a 2.4-approximation in planar graphs [3] and has an $O(\log{n}\log{\log{n}})$-approximation in general graphs [7].
The following is the natural LP for DFVS and its dual, where $\mathcal{C}$ is the set of directed cycles of our graph.
$\min~{}~{}$ $c^{T}x$ ($\mbox{P}_{\text{DFVS}}$) s.t. $x(C)\geq 1\quad\forall\ C\in\mathcal{C}$ (1) $x\geq 0\enspace$ $\operatorname{max}~{}~{}$ $\mathbb{1}^{T}y$ ($\mbox{D}_{DFVS}$) s.t. $\sum_{C\in\mathcal{C},v\in C}y_{C}\leq c_{v}\quad\forall v\in V(G)$ $y\geq\mathbb{0}\enspace.$
Given that constant approximations for DFVS exist for planar graphs, one naturally wonders if DFVS admits constant approximations in bounded genus graphs. We answer this question positively.
Theorem 1.
[…]
From the proof of Theorem 1 it is clear that the algorithm in Theorem 1 runs in time $O(g)\text{poly}(|V(G)|)$.
For uniform costs, the dual LP ($\mbox{D}_{DFVS}$) is the natural LP of the dicycle packing problem. The dicycle packing problem is the problem of finding the maximum number of vertex disjoint dicycles of a graph. Schlomberg et al. [22] show that the LP gap of the natural LP for dicycle packing is at most $\Omega(\frac{1}{g^{2}\log g})$ on any graph of genus $g$. Our result then also implies that the minimum size of a DFVS is at most $O(g^{3}\log g)$ the size of a maximum dicycle packing.
1.1 Our techniques
Informally speaking, for a (directed) graph embedded on a surface where each directed cycle bounds a region homeomorphic to the plane, one can apply the same primal-dual techniques in [12, 3] to obtain a constant factor primal-dual approximation.
In the other case, our algorithm will use the natural LP for DFVS to look for a “separator” [5, 8, 9] $S\subset V$ of cost at most a constant times the optimal DFVS such that $G\backslash S$ is of smaller genus. We obtain a directed cycle $C$, the removal of which results in a surface of one smaller genus. Traversing along the dicycle, we may define a “left” and “right” side of the dicycle. Like in [7], we solve the DFVS LP and use the LP values as distances.
If there is no short path leaving $C$ from the left and entering $C$ from the right and vice versa then there is a small separator $S$ such that each dicycle of $G\backslash S$ either does not use any “left arc” that is, an arc coming in or leaving $C$ from the left, or does not use any “right arc” that is, an arc coming in or leaving $C$ from the right. $G$ with all left (resp. right) arcs deleted is of genus at least one less so inductively we can solve within a constant factor DFVS on $G$ with all left (resp. right) arcs deleted. These two solutions together with $S$ form a DFVS of constant times more than the optimum. If such short paths exist but all starting points of such paths and ending points of such paths are far apart the analysis is similar.
The final case is where there are short paths $P_{1},P_{2}$ leaving $C$ from the left and entering $C$ from the right (or vice versa) and the starting point of $P_{2}$ is close to the endpoint of $P_{1}$. We show that $P_{1}$ is “far” from $P_{2}$ so to speak (we are using directed distances so this is not the same as $P_{2}$ being far from $P_{1}$) and compute a suitable separator. We show that the resulting strongly connected components are of smaller genus. We then combine approximations for different components to give an approximation for the original graph.
The presentation in this paper is focused on demonstrating linear dependence on the genus rather than optimizing the constant in Theorem 1.
2 Preliminaries
In our figures, we will use the representation of the torus by taking the unit square $[0,1]\times[0,1]$ and identifying the two pairs of edges $\{0\}\times[0,1],\ \{1\}\times[0,1]$ and $[0,1]\times\{0\},\ [0,1]\times\{1\}$, that is, the point $(0,p)$ is identified with $(1,p)$ and the point $(q,0)$ is identified with the point $(q,1)$.
Throughout this paper, all surfaces are orientable and smooth and all curves are piece-wise smooth. Distances on surfaces will refer to the geodesic (shortest path on the surface) distance. It is well known (see for instance [21]) that a smooth orientable surface $Q$ is diffeomorphic to the $g$-genus torus for some $g$. We henceforth assume our surface is a $g$-genus torus for some $g$.
Let us call a cycle of $G$, or a closed curve embedded on a surface facial if it bounds a region of the surface homeomorphic to the plane. It is well-known that a cycle is facial if and only if it partitions the surface into 2 connected components. For any set $F\subset V$, define $G^{F}$ to be the residual graph, that is, the subgraph of $G$ induced by those vertices that lie in a dicycle of $G\backslash F$.
3 Solving the case of no facial cycles
We first show the LP gap of the natural LP ($\mbox{P}_{\text{DFVS}}$) for $G$ has integrality gap $O(1)$ in the case $G$ contains no facial cycles. As mentioned previously facial cycles can be tackled in a similar way to that in [12].
Lemma 1.
[…]
Proof.
We prove the statement by induction on the genus $g$. The case $g=0$ was shown in [3]. Suppose the statement is true for $g=g^{\prime}$. Let $Q$ be a surface of genus $g$, Let $G$ be a digraph embedded on $Q$.
First, while the optimal solution $\bar{x}$ to ($\mbox{P}_{\text{DFVS}}$) has a vertex $v$ with value $\bar{x}_{v}\geq\frac{1}{24}$ add v to our temporary hitting set $F$. Formally initialize $F=\emptyset$. Obtain set $F\subset V$ as follows initially set $F=\emptyset$. While the optimal solution $\bar{x}$ to ($\mbox{P}_{\text{DFVS}}$) for $G^{F}$ contains a value $\bar{x}_{v}$ which is $\frac{1}{24}$ or more add $v$ to $F$.
Let $F$ denote the final set obtained. Let $\hat{x}$ be an optimal extreme point solution for the DFVS LP ($\mbox{P}_{\text{DFVS}}$) for $G^{F}$, so $\hat{x}_{v}<\frac{1}{24}\ \ \forall v\in V(G^{F})$. Standard results in iterative rounding, see for instance page 14 of [17], show $F$ has cost at most 24 times the optimal value of our LP.
We now seek to define distances on $G^{F}$. By standard LP theory, $\hat{x}$ has rational coordinates. Let $N\in\mathbb{Z}_{>0}$ be such that $N\hat{x}$ and $\frac{1}{12}N$ are integral, call $N\hat{x}_{v}$, the weight of $v$. For a subgraph $H$ of $G^{F}$ define the weighted distance $d_{\{}\omega,H\}(u,v)$ from $u$ to $v$ the minimum weight of a $u$-$v$ path in $H$. Define $d_{\omega}:=d_{\{\omega,G\}}$ Define $d_{\omega,H}(U,W):=\min_{u\in U,w\in W}\ d_{\omega}(u,w)$. Define $d_{\omega}(U,W)=d_{\omega,G}(U,W)$. Define the weighted distance of a closed walk $P^{\prime}=v_{0},v_{1},..,v_{l}v_{0}$, $\omega(P^{\prime})$ to be $\omega(P^{\prime}):=\sum_{i=0}^{l}N\hat{x}_{v_{i}}$. Since $\hat{x}$ is feasible the following result holds.
Proposition 2.
The weighted distance of any (directed) closed walk $P^{\prime}$ is at least $N$.
Since $\hat{x}$ is optimal, there exists a dicycle $C^{1}:=v_{1},v_{2},..,v_{l^{\prime}}$ such that $\sum_{v\in C^{1}}\hat{x}_{v}=1$. The motivation of our definition of weighted distance comes from [7]. In [7], they also scale the LP values of ($\mbox{P}_{\text{DFVS}}$) so that the resulting values are integer. For any vertex $v$ with $\hat{x}_{v}=0$, they “bypass” the vertex, that is, for each out neighbour $u$ of $v$ and in neighbour $w$ of $v$, they add the edge $wu$ to the graph and when they have done this for all neighbours, they delete $v$ from the graph. For any vertex $v$ with $N\hat{x}_{v}>1`$ they replace $v$ by a “chain” of $N\hat{x}_{v}>1`$ vertices $v_{1}\rightarrow v_{2}\rightarrow..\rightarrow v_{N\hat{x}_{v}}$, that is, for $i=1,2,..,N\hat{x}_{v}-1$, $v_{i}v_{i+1}$ is an edge. $wv_{1}$ and $v_{N\hat{x}_{v}}u$ are edges for each in neighbour $w$ and out neighbour $u$. Call this graph $H$.
For any $W\subset V(H)$ they define “layers” $L_{i}=\{v\in H:\ \ d_{H}(W,v)=i\}$ the nodes at distance $i$ from $W$. They show that the cost of all layers $L_{0},L_{1},...$ is $\sum_{v\in V(G)}N\hat{x}_{v}$. This is very useful for us as we will use this to show that one layer in $L_{1},...L_{m}$ has cost at most $\frac{1}{m}\sum_{v\in V}N\hat{x}_{v}$. However the bypassing and replacing a node with a chain operation of [7] do not preserve the genus of the graph. We instead define the notion of weighted distance $d_{\omega}$. Denote the $i$-th layer from $W$ as $L_{i}:=\{v\in V:i\geq d_{\omega}(W,v)>i-\omega(v)\}$ the set of nodes for which the distance from $W$ to $v$ is at most $i$, but for which the distance plus the weight of $v$ is more than $i$. One can see that $v$ lies in $N\hat{x}_{v}$ different $L_{i}$, which is analogous to how $H$ defined in [7] contains $N\hat{x}_{v}$ copies of $v$ each lying in different layers as well. In particular a node of weight 0 does not lie in any $L_{i}$, which is analogous to how a vertex of weight 0 is bypassed in [7].
Consider the embedding of $C^{1}$ on our surface. Given a subgraph $W$ of $G$ denote by $g(W)$ the subset of our surface occupied by a vertex or edge of $W$. We want to define a “small” neighbourhood around $g(C^{1})$, not containing any vertices outside $C^{1}$, which we divide up into “left” of $g(C^{1})$ and “right” of $g(C^{1})$, which we do using the following propositions. These are slightly informal statements of the exact propositions we require, the precise statements appear in Appendix A.
Proposition 3.
(Informal statement of 12 and 13) Given a closed continuous non-self-intersecting curve $C^{\prime}$ embedded on an orientable surface $Q$, we may partition a small open neighbourhood about $C$ into a “left” $L$ and “right” $R$. For any curve $f:[0,1]\rightarrow Q$ disjoint from $C^{\prime}$ except at $f(1)$ the partition allows us to say that $f$ ”reaches” $C^{\prime}$ from either the left or right.
Proposition 4.
(Informal statement of corollary of 15) Let $C^{\prime}$ be a non-facial closed curve. If a curve $h:[0,1]\rightarrow Q$ “leaves” $C^{\prime}$ at a point $h(0)\in C^{\prime}$ from the left and reaches $C^{\prime}$ at a point $h(1)\in C^{\prime}$ from the right, then $h([0,1])$ together with a subcurve of $C^{\prime}$ from $h(0)$ to $h(1)$ is a non-facial closed curve.
We defer the proofs of 3 and 4 for now. We apply 3 to $g(C^{1})$ let $L,R$ be as in 3 so that each $g(e)$ for $e\in E(G)\backslash E(C^{1})$ is disjoint from at least one of $L,R$ and for each $e\in E(G\backslash C^{1})$, $g(e)$ is disjoint from both $L,R$.
Figure 1: $L$ and $R$ from 3 in yellow and red respectively curve $C^{\prime}$ depicted in black. The curve $h$ leaving $C^{\prime}$ from the left and entering from the right is depicted in dark green. The closed curve formed by $h$ and the subcurve of $C^{\prime}$ between $h(0)$ and $h(1)$ depicted in light green forms a non-facial closed curve.
For each arc $uv$ of $G^{F}$ with exactly one endpoint $v$ on $C^{1}$, $g(uv)$ can be parameterized by a (continuous) curve $f:[0,1]\rightarrow g(uv)$ with $f(0)=g(u)$, $f(1)=g(v)$. If $f$ reaches $g(C^{1})$ from the left we say that $uv$ reaches $C^{1}$ from the left, otherwise, we say $uv$ reaches $C^{1}$ from the right.
Let $u^{\prime}_{i,1},u^{\prime}_{i,2},..,u^{\prime}_{i,l_{i}}$ be the out neighbours of $v_{i}$ such that the edges $u^{\prime}_{i,t^{\prime}}v_{i}$ reach $v_{i}$ from the left, that is, the arc obtained from reversing the arc $v_{i}u^{\prime}_{i,t^{\prime}}$ of our graph reaches $v_{i}$ from the left. Let $w^{\prime}_{i,1},w^{\prime}_{i,2},..,w^{\prime}_{i,z_{i}}$ the in neighbours of $v_{i}$ such that the edges $w^{\prime}_{i,t^{\prime}}v_{i}$ reach $v_{i}$ from the right. Subdivide each edge $v_{i}u^{\prime}_{i,t}$ into a path $v_{i}u_{i,t}u^{\prime}_{i,t}$ and each edge $w^{\prime}_{j,t^{\prime}}v_{j}$ into a path $w^{\prime}_{j,t^{\prime}}w_{j,t^{\prime}}v_{j}$ and give the new vertices $w_{j,t^{\prime}},u_{i,t}$ infinite cost. There is a natural embedding of our new graph on our surface by placing each $u_{i,t}$ where the midpoint of the curve $g(v_{i}u^{\prime}_{i,t})$ was embedded and likewise for $w_{j,t^{\prime}}$. By abuse of notation, we continue to call our graph $G$ and define $\hat{x}_{u_{i,t}}=\hat{x}_{w_{j,t^{\prime}}}=0$ for all $u_{i,t}$, $w_{j,t^{\prime}}$. Denote $U:=\cup_{i=1}^{l^{\prime}}\{u_{i,1},u_{i,2},..,u_{i,l_{i}}\}$ and $W:=\cup_{i=1}^{l^{\prime}}\{w_{i,1},w_{i,2},..,u_{i,l_{i}}\}$. For $X\subset[l^{\prime}]$, denote $U_{X}:=\cup_{i\in X}\{u_{i,1},u_{i,2},..,u_{i,l_{i}}\}$, $V_{X}=\{v_{i}:\ \ i\in X\}$ and $W_{X}:=\cup_{i\in X}\{w_{i,1},w_{i,2},..,w_{i,z_{i}}\}$.
Let $\tau_{-}:=\{i\in[l^{\prime}]:\exists w_{i,t^{\prime}}\in W,\ \ \exists u_{j,t}\in U:d_{\omega,G^{F}\backslash C^{1}}(u_{j,t},w_{i,t^{\prime}})<\frac{1}{12}N\}$ the first indices of the set of vertices of $W$ of weighted distance at most $\frac{1}{12}N$ from $U$ in $G^{F}\backslash C^{1}$. Let $\tau_{+}:=\{j\in[l^{\prime}]:\exists u_{j,t}\in U,\ \ \exists w_{i,t^{\prime}}\in W:d_{\omega,G^{F}\backslash C^{1}}(u_{j,t},w_{i,t^{\prime}})<\frac{1}{12}N\}$ the first indices of the set of vertices of $U$ that can reach $W$ in $G^{F}\backslash C^{1}$ with a path of weighted distance at most $\frac{1}{12}N$.
Claim 1.
If $d_{\omega}(V_{\tau_{-}},V_{\tau_{+}})>\frac{1}{12}N$, then we can find $S\subset V$, $c(S)=O(1)OPT_{LP}$, where $OPT_{LP}:=\sum_{v\in V}c_{v}x_{v}$ is the value of the optimal fractional solution, such that any strongly connected component of $G^{F}\backslash S$ does not contain a directed path from $U$ to $W$ in $G\backslash C^{1}$.
[…]
Proof.
Suppose $d_{\omega,G\backslash C^{1}}(V_{\tau_{-}},V_{\tau_{+}})>\frac{1}{12}N$. For $i=0,..,\frac{1}{12}N$ let $S_{i}:=\{v\in V:i\geq d_{\omega,G^{F}\backslash C^{1}}(U\backslash U_{\tau_{+}},v)>i-\omega(v)\}$ denote the set of vertices of $V$ that are at weighted distance $i$ from $U\backslash U_{\tau_{+}}$ in $G^{F}\backslash C^{1}$. (see Figure 2). Since $d_{\omega,G^{F}\backslash C^{1}}(U\backslash U_{\tau_{+}},W)>\frac{1}{12}N$, for $i=0,..,\frac{1}{12}N$, $W\cap S_{i}=\emptyset$ and $W$ is not reachable from $U\backslash U_{\tau_{+}}$ in $(G^{F}\backslash C^{1})\backslash S_{i}$ for any $i$.
Figure 2: Nodes of $U_{\tau_{+}}$ and $V_{\tau_{-}}$ shown in blue.
Since each $v$ can lie in at most $\omega(v)$ $S_{i}$, $\sum_{i=0}^{\frac{1}{12}N}c(S_{i})\leq N\cdot OPT_{LP}$. Let $S^{\prime}$ be the $S_{i}$ of minimum cost. For $i=0,..,\frac{1}{12}N$ let $T_{i}:=\{v\in V:i>d_{\omega,G^{F}\backslash C^{1}}(v,W\backslash W_{\tau_{-}})-\omega(v),\ \ d_{\omega,G^{F}\backslash C^{1}}(v,W\backslash W_{\tau_{-}})\geq i\}$. Since for $v\in U$, $d_{\omega,G^{F}\backslash C^{1}}(v,W\backslash W_{\tau_{-}})-\omega(v)=d_{\omega}(v,W\backslash W_{\tau_{-}})>\frac{1}{12}N$, $U\cap T_{i}=\emptyset$ for $i=0,..,\frac{1}{12}N$. Hence $W\backslash W_{\tau_{-}}$ is not reachable from $U$ in $(G^{F}\backslash C^{1})\backslash T_{i}$ for any $i$. Let $T^{\prime}$ be the $T_{i}$ of minimum cost.
Finally, let $Y_{i}:=\{v\in V:\ i\geq d_{\omega}(V_{\tau_{-}},v)>i-\omega(v)\}$ the set of vertices of weighted distance $i$ from $V_{\tau_{-}}$. By assumption $d_{\omega}(V_{\tau_{-}},V_{\tau_{+}})>\frac{1}{12}N$ and hence $V_{\tau_{+}}$ is not reachable from $V_{\tau_{-}}$ in $G^{F}\backslash Y_{i}$ for any $i=1,2,..,\frac{1}{12}N$. Let $Y^{\prime}$ be the $Y_{i}$ of minimum cost.
Let $S:=S^{\prime}\cup T^{\prime}\cup Y^{\prime}$. We claim no strongly connected component $K^{\prime}$ of $G^{F}\backslash S$ contains a directed path from $U$ to $W$ in $G^{F}\backslash C^{1}$. Suppose for a contradiction that some strongly connected component $K^{\prime}$ of $G^{F}\backslash S$ contains a directed path from some $u_{i,t}\in U$ to some $w_{j,t^{\prime}}\in W$.
If $j\notin\tau_{-}$, then $w_{j,t^{\prime}}$ is not reachable from $U$ in $G^{F}\backslash S$. If $i\notin\tau_{+}$, then $W$ is not reachable from $u_{i,t}$ in $G^{F}\backslash S$. Thus if either $j\notin\tau_{-}$ or $i\notin\tau_{+}$ then there is no path from $u_{i,t}$ to $w_{j,t^{\prime}}$ in $G^{F}\backslash S$. Thus, $j\in\tau_{-}$ and $i\in\tau_{+}$. As $K^{\prime}$ is strongly connected, this implies that $G^{F}\backslash S$ contains a path from $V_{\tau_{-}}$ to $V_{\tau_{+}}$, which is not possible.
Now suppose that $d_{\omega}(V_{\tau_{-}},V_{\tau_{+}})\leq\frac{1}{12}N$. Let $i\in\tau_{-}$ and $j\in\tau_{+}$ be such that $d_{\omega}(v_{i},v_{j})\leq\frac{1}{12}N$ Let $P_{1},P_{2},P_{3}$ be $u_{a,t}$-$v_{i}$, $u_{j,t^{\prime}}$-$v_{b}$ and $v_{i}$-$v_{j}$ paths of weight at most $\frac{1}{12}N$, with the second last vertices of $P_{1},P_{2}$ being in $W$, for some $a,b$. Such paths exist as $i\in\tau_{-}$ and $j\in\tau_{+}$. If $a=i$, then $P_{1}v_{i}u_{a,t}$ is a cycle for which $\sum_{v\in P_{1}v_{i}u_{a,t}}\hat{x}_{v}<1$ which is a contradiction. So $a\neq i$, likewise $b\neq j$.
For $i^{\prime},j^{\prime}\in\{1,2,..,l^{\prime}\}$, let $C^{1}_{(v_{i^{\prime}},v_{j^{\prime}})}:=v_{i^{\prime}},v_{i^{\prime}+1},v_{i^{\prime}+2},..,v_{j^{\prime}-1}v_{j^{\prime}}$ (where $v_{t}=v_{t\pmod{l^{\prime}}}$) denote the directed path in $C^{1}$ from $v_{i}$ to $v_{j}$. Note that $d_{\omega}(v_{i^{\prime}},v_{j^{\prime}})=\omega(C^{1}_{(v_{i^{\prime}},v_{j^{\prime}})})$, for otherwise there is a $v_{i^{\prime}}$-$v_{j^{\prime}}$ path $P^{\prime}$ of weight less than $d_{\omega}(v_{i^{\prime}},v_{j^{\prime}})$. Then $C^{1}_{(v_{j^{\prime}},v_{i^{\prime}})}\cup P^{\prime}$ is a directed closed walk of weight $\omega(C^{1})-\omega(C^{1}_{(v_{i^{\prime}},v_{j^{\prime}})})+d_{\omega}(v_{i^{\prime}},v_{j^{\prime}})<\omega(C^{1})$. Noting that the weighted distance of a cycle is equal to $N\sum_{v\in C^{1}}\hat{x}_{v}$, we obtain $N\sum_{v\in C^{1}_{(v_{j^{\prime}},v_{i^{\prime}})}\cup P^{\prime}}\hat{x}_{v}<\omega(C^{1})=N$, from which it follows the sum of the $\hat{x}_{v}$ values along the closed walk $C^{1}_{(v_{j^{\prime}},v_{i^{\prime}})}\cup P^{\prime}$, $\sum_{v\in C^{1}_{(v_{j^{\prime}},v_{i^{\prime}})}\cup P^{\prime}}\hat{x}_{v}$ is strictly less than 1, which contradicts the feasibility of $\hat{x}$.
We claim $\omega(C^{1}_{(v_{a},v_{i})}),\omega(C^{1}_{(v_{j},v_{b})})\leq\frac{1}{12}N$. Suppose for a contradiction that $\omega(C^{1}_{(v_{a},v_{i})})>\frac{1}{12}N$ Since $\omega(C^{1})=N$, this implies that $\omega(C^{1}_{(v_{i+1},v_{a-1})})<N-\frac{1}{12}N$. Then the cycle $P_{1}C^{1}_{(v_{i},v_{a})}v_{a}u_{a,t}$ satisfies $\sum_{v\in V(P_{1}C^{1}_{(v_{i},v_{a})}v_{a}u_{a,t})}\hat{x}_{v}<1-\frac{1}{12}+\frac{1}{12}=1$ which is a contradiction. Likewise $\omega(v_{j},v_{b})\leq\frac{1}{12}N$.
Let us show $C^{1}_{(v_{a},v_{i})}\cap C^{1}_{(v_{j},v_{b})}=\{v_{i}\}\cap\{v_{j}\}$, that is the paths $C^{1}_{(v_{a},v_{i})}$ and $C^{1}_{(v_{j},v_{b})}$ are disjoint except in the case $i=j$ when their intersection is $v_{i}$. First, let us address the case $i\neq j$. Suppose for a contradiction that $C^{1}_{(v_{a},v_{i})}\cap C^{1}_{(v_{j},v_{b})}\neq\emptyset$. Let $v\in C^{1}_{(v_{a},v_{i})}\cap C^{1}_{(v_{j},v_{b})}$. Note $v\neq v_{i},v_{j}$ for otherwise $C^{1}_{(v_{j},v_{i})}P_{3}$ is a closed walk of weight less than $N$. Let $Q_{1}$ be a path from $v$ to $v_{i}$ in $C^{1}_{(v_{a},v_{i})}$ and $Q_{2}$ a path from $v_{j}$ to $v$ in $C^{1}_{(v_{j},v_{b})}$. Then $Q_{2}Q_{1}P_{3}$ is a closed walk of weighted distance at most $\frac{1}{4}N$ which is a contradiction.
Now suppose that $i=j$. Suppose for a contradiction that $C^{1}_{(v_{a},v_{i})}\cap C^{1}_{(v_{i},v_{b})}\neq\{v_{i}\}$. Let $v\in(C^{1}_{(v_{a},v_{i})}\cap C^{1}_{(v_{i},v_{b})})\backslash v_{i}$. Let $Q_{1}$ be a path from $v$ to $v_{i}$ in $C^{1}_{(v_{a},v_{i})}$ and $Q_{2}$ a path from $v_{i}$ to $v$ in $C^{1}_{(v_{i},v_{b})}$. Then $Q_{1}Q_{2}$ is a closed walk of weighted distance at most $\frac{1}{6}N$, which is a contradiction.
Claim 2.
$d_{\omega}(P_{2}\cup C^{1}_{(v_{j+1},v_{b})},P_{1}\cup C^{1}_{(v_{a},v_{i})})\geq\frac{1}{6}N$.
Proof.
Suppose for a contradiction that $d_{\omega}(P_{2}\cup C^{1}_{(v_{j+1},v_{b})},P_{1}\cup C^{1}_{(v_{a},v_{i})})<\frac{1}{6}N$. Let $s\in P_{2}\cup C^{1}_{(v_{j+1},v_{b})}$ and $q\in P_{1}\cup C^{1}_{(v_{a},v_{i})}$ be such that $d_{\omega}(s,q)<\frac{1}{6}N$. Let $P^{\prime}_{1}$ be the directed path in $P_{1}\cup C^{1}_{(v_{a},v_{i})}$ from $q$ to $v_{i}$. $P^{\prime}_{2}$ the directed path in $P_{2}\cup C^{1}_{(v_{j},v_{b})}$ from $v_{j}$ to $s$ and $Q$ a path of weight at most $\frac{1}{6}N$ from $s$ to $q$. Then $\bar{C}:=v_{j}P^{\prime}_{2}QP^{\prime}_{1}P_{3}$ is a closed walk such that $\sum_{v\in\bar{C}}\hat{x}_{v}<1$ which is a contradiction (see Figure 3).
Thus, $d_{\omega}(P_{2}\cup C^{1}_{(v_{i},v_{b})},P_{1}\cup C^{1}_{(v_{a},v_{i})})\geq\frac{1}{6}N$. Since $d_{\omega}(u,v_{i})\leq\frac{1}{12}N$ for any $u\in C^{1}_{(v_{a},v_{i})}$, $d_{\omega}(P_{2},C^{1}_{(v_{a},v_{i})})\geq\frac{1}{12}N$
For $i=0,1,..,\frac{1}{12}N$, define $R_{i}:=\{v\in V:i\geq d_{\omega}(P_{2}\cup C^{1}_{(v_{j+1},v_{b})},v)>i-\omega(v)\}$. Each vertex $v\in V$ lies in at most $\omega(v)$ $R_{i}$. Let $R^{\prime}$ be the $R_{i}$ of the smallest cost, so $c(R^{\prime})\leq{12}OPT_{LP}$. Since $d_{\omega}(P_{2}\cup C^{1}_{(v_{j+1},v_{b})},P_{1}\cup C^{1}_{(v_{a},v_{i})})\geq\frac{1}{12}N$, it follows that $(P_{1}\cup C^{1}_{(v_{a},v_{i})})\backslash R_{i}$ is not reachable from $(P_{2}\cup C^{1}_{(v_{j+1},v_{b})})\backslash R_{i}$ in $G\backslash R_{i}$ for any $i$. Thus, $(P_{1}\cup C^{1}_{(v_{a},v_{i})})\backslash R^{\prime}$ is not reachable from $(P_{2}\cup C^{1}_{(v_{j+1},v_{b})})\backslash R^{\prime}$ in $G\backslash R^{\prime}$.
Thus, any strongly connected component of $G\backslash R^{\prime}$ is either contained in $G^{F}\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or $G^{F}\backslash(P_{2}\cup C^{1}_{(v_{j+1},v_{b})})$. For $i=1,2,..,\frac{1}{12}N$ let $K^{+}_{i}:=\{v\in V:\ \ i\geq d_{\omega}(v_{j},v)>i-\omega(v)\}$ be the vertices of weighted distance $i$ from $v_{j}$. Let $K^{\prime+}$ denote the $K^{+}_{i}$ of minimum cost.
Figure 3: On the left, there are $u_{1,1}$-${v_{2}}$ and $u_{2,1}$-${v_{3}}$ paths (green and blue vertices respectively) of weight at most $\frac{1}{12}N$ and $s$-$q$ path of length at most $\frac{1}{12}N$. The red cycle would then have weight at most $N$, which is a contradiction. On the right are the sets $R_{i}$, vertices at distance $i$ from $P_{2}\cup C^{1}_{(v_{j+1},v_{b})}$.
Claim 3.
$v_{j}$ is not contained in a cycle in $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$.
Proof.
Suppose that there is a cycle $a_{1},a_{2},..,a_{p}v_{j}a_{1}$ in $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$. If $d_{\omega}(v_{j},a_{p})>\frac{1}{12}N$, then $a_{p}$ is not reachable from $v_{j}$ in $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$. Thus, there is a path $P_{a}$ of weighted distance at most $\frac{1}{12}N$ from $v_{j}$ to $a_{p}$. Then the closed walk $v_{j}p_{a}a_{p}v_{j}$ has weighted distance at most $\omega(P_{a})+\omega(a_{p})\leq\frac{1}{12}N+\frac{1}{12}N<N$ which is a contradiction. ∎
Recall that any dicycle of $G^{F}\backslash R^{\prime}$ is contained in either $G^{F}\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or $G^{F}\backslash(P_{2}\cup C^{1}_{(v_{j+1},v_{b})})$. Since $v_{j}$ is not contained in any dicycle of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$ it follows that any dicycle of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$ is either contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or in $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$. By 4, $g(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ and $P_{2}\cup C^{1}_{(v_{j},v_{b})}$ are nonfacial.
Definition 1.
[1, 13] Given a simple closed curve $f$ on a surface without boundary $Q$, not dividing the surface into 2 regions, we say $Q^{\prime}$ is obtained by doing surgery along $f$ if $Q^{\prime}$ is obtained as follows. “Thicken” $f$ to obtain a cylinder and remove this cylinder from $Q$, call this resulting surface $Q^{\prime\prime}$. The boundary of $Q^{\prime\prime}$ consists of 2 circles we “glue” two cones along these circles and call this final surface $Q^{\prime}$.
Theorem 5.
([1] p.162) For a surface without boundary $Q$ of genus $g^{\prime}$, $Q^{\prime}$ obtained by 1 is a surface without boundary of genus at most $g^{\prime}-1$.
We apply the surgery of 1 to $g(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ to obtain a surface $Q^{\prime}$ of genus one less than $Q$. Thus, $G^{F}\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ is of genus at least one less than $G^{F}$. Likewise $G^{F}\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$ is of genus at least one less than $G^{F}$. Now let $G_{1},G_{2},..,G_{l}$ be the strongly connected components of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$. Since any closed walk of of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$ is either contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or in $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$ each strongly connected component is contained in either contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or in $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$. Since both $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ and $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$ are of smaller genus, each $G_{i}$ is of smaller genus. By induction, there are solutions $A_{i}$ to $G_{i}$ of cost $c_{g-1}OPT_{LP(G_{i})}$, where $c_{g}$ is the integrality gap of the DFVS LP for graphs of genus $g$.
Define $\hat{x}^{G}_{i}\in\mathbb{R}^{V(G_{i})}$ as $\hat{x}^{G}_{i}(v)=\hat{x}_{v}$, where $\hat{x}$ is as in the proof of 1. Define $\hat{x}^{G}_{i}\in\mathbb{R}^{V(G_{i})}$ as $\hat{x}^{G}_{i}(v)=\hat{x}_{v}$, where $\hat{x}$ is as in the proof of 1. Since graphs $G_{i}$ are vertex disjoint, $\sum_{i=1}^{l}OPT_{LP(G_{i})}\leq\sum_{i=1}^{l}\sum_{v\in V(G_{i})}\hat{x}^{G}_{i}(v)\leq\sum_{v\in V(G)}\hat{x}=OPT_{LP(G)}$. Now $F\cup R^{\prime}\cup K^{\prime+}\cup(\cup_{i=1}^{l}A_{i})$ is a DFVS of cost $(O(1)+c_{g-1})OPT_{LP(G)}=(O(1)+O(g-1))OPT_{LP(G)}=(O(g))OPT_{LP(G)}$.
Let $b^{\prime}_{i,1},b^{\prime}_{i,2},..,b^{\prime}_{i,l^{\prime}_{i}}$ be the in neighbours of $v_{i}$ such that each edge $b^{\prime}_{i,t}v_{i}$ reaches $v_{i}$ from the left and $d^{\prime}_{i,1},d^{\prime}_{i,2},..,d^{\prime}_{i,t^{\prime}_{i}}$ be the out neighbours of $v_{i}$ such that the edge $d^{\prime}_{i,t^{\prime}}v_{i}$ reaches $v_{i}$ from the right. Subdivide each edge $b^{\prime}_{j,t^{\prime}}v_{j}$ into a path $b^{\prime}_{j,t^{\prime}}b_{j,t^{\prime}}v_{j}$ and each edge $v_{i}d^{\prime}_{i,t}$ into a path $v_{i}d_{i,t}d^{\prime}_{i,t}$ and give the new vertices $d_{j,t^{\prime}},b_{i,t}$ infinite cost. There is a natural embedding of our new graph on our surface by placing each $b_{i,t}$ where the midpoint of the curve $g(v_{i}b^{\prime}_{j,t^{\prime}})$ was embedded and likewise for $d_{j,t^{\prime}}$. By abuse of notation, we continue to call our graph $G$ and define $\hat{x}_{b_{i,t}}=\hat{x}_{d_{j,t^{\prime}}}=0$ for all $b_{i,t}$, $d_{j,t^{\prime}}$.
Denote $B:=\cup_{i=1}^{l^{\prime}}\{b_{i,1},b_{i,2},..,b_{i,l^{\prime}_{i}}\}$, $D=\cup_{i=1}^{l^{\prime}}\{d^{\prime}_{i,1},d^{\prime}_{i,2},..,d^{\prime}_{i,t^{\prime}_{i}}\}$. Let $\kappa_{-}:\{i\in[l^{\prime}]:\exists b_{i,t^{\prime}}\in D,\ \ \exists d_{j,t}\in U:d_{\omega,G^{F}\backslash C^{1}}(d_{j,t},b_{i,t^{\prime}})<\frac{1}{12}N\}$ the first indices of the set of vertices of $B$ of weighted distance at most $\frac{1}{12}N$ from $D$ in $G^{F}\backslash C^{1}$. Let $\kappa_{+}:=\{j\in[l^{\prime}]:\exists d_{j,t}\in U,\ \ \exists b_{i,t^{\prime}}\in W:d_{\omega,G^{F}\backslash C^{1}}(d_{j,t},b_{i,t^{\prime}})<\frac{1}{12}N\}$ the first indices of the set of vertices of $D$ that can reach $B$ with a path of weighted distance at most $\frac{1}{12}N$ in $G^{F}\backslash C^{1}$. Similarly to how we proved 1, we can show the following:
Claim 4.
If $d_{\omega}(V_{\kappa_{-}},V_{\kappa_{+}})>\frac{1}{12}N$, then we can find $T\subset V$, $c(T)=O(1)OPT_{LP}$, (recall $OPT_{LP}:=\sum_{v\in V}c_{v}x_{v}$ is the value of the optimal fractional solution), such that any strongly connected component of $G^{F}\backslash T$ does not contain a directed path from $D$ to $B$ in $G\backslash C^{1}$.
[…]
We now construct a DFVS of cost at most $O(g)OPT_{LP}$. If either $d_{\omega}(V_{\kappa_{-}},V_{\kappa_{+}})\leq\frac{1}{12}N$ or $d_{\omega}(V_{\tau_{-}},V_{\tau_{+}})\leq\frac{1}{12}N$. Then 1 or 4 respectively shows that that the LP gap of the natural LP ($\mbox{P}_{\text{DFVS}}$) for $G$ has integrality gap $O(g)$.
Now assume both $d_{\omega}(V_{\tau_{-}},V_{\tau_{+}}),d_{\omega}(V_{\kappa_{-}},V_{\kappa_{+}})>\frac{1}{12}N$. Then by 1 and 4, there are sets $S,T$ such that any strongly connected component of $G^{F}\backslash(S\cup T)$ does not contain a path from $U$ to $W$ or a path from $D$ to $B$ in $G^{F}\backslash C^{1}$.
For any digraph $H$ define $\operatorname{un}(H)$ to be the underlying (undirected) graph of $H$. Let $K$ be any strongly connected component of $G^{F}\backslash(S\cup T)$. We will prove $\operatorname{un}(K)$ does not contain any path from $U\cup B$ to $W\cup D$ in $\operatorname{un}(K)\backslash C^{1}$.
Proposition 6.
If there is a (undirected) path $P=u_{i,t}q_{1},q_{2},..,q_{t}$ from some $u_{i,t}\in U$ (resp $u_{i,t}\in D$) in $\operatorname{un}(K)\backslash C^{1}$, then there is a directed path from $U$ (resp $D$) to $q_{j}$ in $G^{F}\backslash(S\cup T\cup C^{1})$ for any $j=1,2,..,t$.
If there is a (undirected) path $P=q_{1},q_{2},..,q_{t}w_{i,t}$ from some $w_{i,t}\in W$ (resp $b_{i,t}\in B$) in $\operatorname{un}(K)\backslash C^{1}$, then there is a directed path from $U$ (resp to $q_{j}$ in $G^{F}\backslash(S\cup T\cup C^{1})$ for any $j=1,2,..,t$.
Proof.
Let $P=u_{i,t}q_{1},q_{2},..,q_{t}$ be a path in $\operatorname{un}(K)\backslash C^{1}$ from some $u_{i,t}\in U$ (resp $u_{i,t}\in D$). We prove by induction $t^{\prime}$ on that there is a directed path from $U$ to $q_{j}$ in $G^{F}\backslash(S\cup T\cup C^{1})$ for any $j=1,2,..,t^{\prime}$. The case $t^{\prime}=1$ is clear as each $u_{i,t}\in U$ (resp $u_{i,t}\in D$) only has out neighbours so the undirected edge $\{u_{i,t},q_{1}\}$ in $\operatorname{un}(K)$ is directed from $u_{i,t}$ to $q_{1}$.
Now assume the statement true for $t^{\prime}=t^{\prime\prime}$. For $t^{\prime}=t^{\prime\prime}+1$, if the undirected edge $\{q_{t^{\prime}},q_{t^{\prime}+1}\}$ is directed from $q_{t^{\prime}}$ to $q_{t^{\prime}+1}$, then there is a directed path from $u_{i,t}$ to $q_{j}$ in $G^{F}\backslash(S\cup T\cup C^{1})$.
Otherwise $\{q_{t^{\prime}},q_{t^{\prime}+1}\}$ is directed from $q_{t^{\prime}+1}$ to $q_{t^{\prime}}$. By strong connectedness of $K$, there is a directed path $P^{\prime}$ from $q_{t^{\prime}}$ to $q_{t^{\prime}+1}$ in $K\backslash(S\cup T)$. $P^{\prime}$ must intersect $C^{1}$. Thus $P^{\prime}$ must intersect $W$ or $B$. Let $P^{\prime\prime}$ denote the subpath of $P^{\prime}$ from $u_{i,t}$ to when $P^{\prime}$ first intersects $U$ or $B$. By construction $P^{\prime\prime}$ lies in $G^{F}\backslash(S\cup T\cup C^{1})$. As $u_{i,t}$ lies in $U$ (resp $D$) $P^{\prime\prime}$ does not intersect $W$ (resp. $B$). Thus $P^{\prime\prime}$ is a $u_{i,t}$-$B$ (resp. $u_{i,t}$-$W$) path. Consider the subpath $Q$ of the reversal of $P^{\prime}$ starting from $q_{t^{\prime}+1}$ to when the reversal of $P^{\prime}$ first intersects $D$ or $U$. Let $\operatorname{rev}(Q)$ denote the reversal of $Q$. Note $\operatorname{rev}(Q)$ lies in $G^{F}\backslash(S\cup T\cup C^{1})$. If the starting vertex of $\operatorname{rev}(Q)$ is in $D$ (resp. $U$), then $\operatorname{rev}(Q)\cup P^{\prime}$ is a $D$-$B$ (resp. $U$-$W$) path in $G^{F}\backslash(S\cup T\cup C^{1})$. This contradicts 4. Thus the starting vertex of of $\operatorname{rev}(Q)$ is in $U$ (resp. $D$). This implies there is a path from $U$ (resp. $D$ ) to $q_{t^{\prime}}$ completing the induction. The proof of the second part is similar. ∎
Proposition 7.
There is no (undirected) path from $W\cup D$ to $U\cup B$ in $\operatorname{un}(K)\backslash C^{1}$.
Proof.
If we have a $U$-$W$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}w_{j,t^{\prime}}$ in $\operatorname{un}(K)\backslash C^{1}$, then by 6, there are directed $U$-$q_{1}$ and $q_{1}$-$W$ paths $P_{1}$ and $P_{2}$ in $\operatorname{un}(K)\backslash C^{1}$. Then $P_{1}\cup P_{2}$ is a directed $U$-$W$ path in $K\backslash C^{1}$ which contradicts 4. Thus we do not have a $U$-$W$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}w_{j,t^{\prime}}$ in $K\backslash C^{1}$. Likewise we do not have a $D$-$B$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}w_{j,t^{\prime}}$ in $K\backslash C^{1}$.
Suppose we have a $U$-$D$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}w_{j,t^{\prime}}$ in $\operatorname{un}(K)\backslash C^{1}$. By 6, there are directed $U$-$q_{1}$ and $D$-$q_{1}$ paths $P_{1}$ and $P_{2}$ in $K\backslash C^{1}$. Recall $U$ has no in-neighbours of in $G\backslash C^{1}$, so the edge $\{u_{i,t},q_{1}\}$ in $K$ is directed from $u_{i,t}$ to $q_{1}$. By 2 connectedness of $K$, there is a path $P_{3}$ from $q_{1}$ to $u_{i,t}$. The only in-neighbours of $u_{i,t}$ are in $C^{1}$, thus $P_{3}$ intersects $W\cup B$. Let $P^{\prime}_{3}$ be the subpath of $P_{3}$ from $q_{1}$ to when it the path first intersects $W\cup B$. If the endpoint of $P^{\prime}_{3}$ is in $W$, then $P_{1}\cup P^{\prime}_{3}$ is a $U$-$W$ path in $K\backslash C^{1}$. Otherwise, if the endpoint of $P^{\prime}_{3}$ is in $B$, then $P_{2}\cup P^{\prime}_{3}$ is a $D$-$B$ path in $K\backslash C^{1}$. Either way this contradicts 4.
Proposition 8.
[14][24]Suppose $G$ is a graph embedded on a surface $Q$. Let $C$ be a cycle of $G$ that does not divide $Q$ into two separate regions such that there is no edge between vertices of $C$ that is not part of $C$. Define a “left and right as in [] Let $\hat{L}$, $\hat{R}$ denote the neighbours of $C$ that are “left” or “right” of $C$. Suppose each connected component of $G\backslash C$ only contains nodes of $\hat{L}$ or $\hat{R}$ but not both. Then $G$ is embeddable on a surface of smaller genus.
Applying 7, we get that each strongly connected component $K$ of $G^{F}\backslash C^{1}$ satisfies 8. Thus $K$ is embeddable on a surface of $g-1$. Let $G_{1},G_{2},.,G_{l}$ be the strongly connected components of $G^{F}\backslash(S\cup T)$, so each $G_{i}$ is embeddable on a surface of $g-1$. By induction the LP gap for any strongly connected component $K$ of $G^{F}\backslash C^{1}$ is at most some constant $c_{g-1}$. By induction, there are solutions $A_{i}$ to $G_{i}$ of cost $c_{g-1}OPT_{LP(G_{i})}$, where $c_{g}$ is the integrality gap of the DFVS LP for graphs of genus $g$. Define $\hat{x}^{G}_{i}\in\mathbb{R}^{V(G_{i})}$ as $\hat{x}^{G}_{i}(v)=\hat{x}_{v}$, where $\hat{x}$ is as in the proof of 1. Since graphs $G_{i}$ are vertex disjoint, $\sum_{i=1}^{l}OPT_{LP(G_{i})}\leq\sum_{i=1}^{l}\sum_{v\in V(G_{i})}\hat{x}^{G}_{i}(v)\leq\sum_{v\in V(G)}\hat{x}=OPT_{LP(G)}$. Then $S\cup T\cup F\cup(\cup_{i=1}^{l}A_{i})$ is a DFVS of cost $(O(1)+c_{g-1})OPT_{LP(G)}=(O(1)+O(g-1))OPT_{LP(G)}=(O(g))OPT_{LP(G)}$.
As observed in [7], ($\mbox{P}_{\text{DFVS}}$) can be solved in polynomial-time via the ellipsoid method. Hence 1 yields a polynomial time $O(g)$-approximation algorithm for DFVS in graphs of genus $g$ with no facial cycle.
4 Solving the general case
We now show how to obtain an $O(g)$-approximation for DFVS for graphs embedded on any surface $Q$ of genus $g$.
See 1
Proof.
We prove the statement by induction on the genus $g$. The case $g=0$ was shown in [3]. Suppose the statement is true for $g=g^{\prime}$. Let $Q$ be a surface of genus $g$. Let $G$ be a digraph embedded on $Q$. If $C$ is a facial cycle bounding a face of $G$ call $C$ face-minimal. Note that if $G$ contains a facial cycle, then it must contain a face minimal cycle by the following argument. We associate each facial cycle $C$ with the region $R$ of $Q$ homeomorphic to the plane that it bounds. Let $C$ be a dicycle bounding a minimal (by containment) region of $Q$. If $C$ is not a face then there is a dipath $P$ between two nodes $u,v$ of $C$ then $P$ together with either the $u$-$v$ or $v$-$u$ dipath in $C$ forms a cycle bounding a smaller region of $Q$, which is a contradiction.
Our algorithm proceeds as follows. The first part is a primal-dual algorithm analogous to the technique of [12] for DFVS in planar graphs. Given feasible dual solution $y$ to ($\mbox{D}_{DFVS}$), let the residual cost of node $v\in V$ be $c_{v}-\sum_{C\in\mathcal{C},v\in C}y_{C}$. For $\hat{S}\subset V(G)$, recall $G^{\hat{S}}$ denotes the subgraph of $G$ induced by those vertices which are in a dicycle of $G\backslash\hat{S}$.
Our primal-dual method begins with a trivial feasible dual solution $y=\mathbb{0}$, and the empty, infeasible hitting set $\hat{S}=\emptyset$.
While $G^{\hat{S}}$ contains a facial cycle, increment the dual variables $y_{C}$ in $\mbox{P}_{\text{DFVS}}$ of face minimal cycles $C$ of $G$. When a node of $G$ becomes tight add it to $\hat{S}$. When $G^{\hat{S}}$ contains no facial cycles apply reverse deletion to $\hat{S}$ with respect to the facial cycles of $G^{\hat{S}}$, that is, we consider each node $v$ of $\hat{S}$ in the order it was added and if $G\backslash(\hat{S}\backslash\{v\})$ contains no facial cycles, delete $v$ from $\hat{S}$. Denote by $\bar{S}$ the set $\hat{S}$ at the end of the algorithm. In other words, we apply the primal-dual method to solve the problem of hitting all facial dicycles of $G$.
In the second step using 1, we find a DFVS $\tilde{S}$ of $G^{\bar{S}}$ of cost at most $O(g)OPT_{LP}$.
Input : A digraph $G=(V,E)$ with non-negative node-costs $c_{v}$, for each $v\in V$.
Output : A Directed FVS $S$ of $G$.
1 $S=\emptyset$
2 while $G^{S}$ contains a facial cycle do
3       Increment all dual variables $y_{C}$ for face minimal cycles of $G^{S}$. Add all nodes that became tight to $S$.
4 end while
5Reverse-Deletion:
6       Let $s_{1},s_{2},..,s_{l}$ be nodes of $S$ in the order they were added.
7       for  $t=l$ downto $1$ do
8             if  $G^{S\backslash\{s_{t}\}}$ contains no facial cycle  then
9                  $S\leftarrow S\backslash\{s_{t}\}$
10             end if
11
12       end for
13
14
15
16Use 1 to find FVS $\tilde{S}$ of $G^{\bar{S}}$ of cost at most $O(g)OPT_{LP}$
return $S\cup\tilde{S}$
Algorithm 1 MinWeightDirectedFVS $(G,c)$
Clearly, $\bar{S}\cup\tilde{S}$ is a feasible FVS, we claim it has cost $O(g)OPT_{LP}$. It suffices to prove that $\bar{S}=O(g)OPT_{LP}$. To do so we apply that standard analysis of primal-dual methods in [11, 12].
Theorem 9.
[…]
  1. 1.
    $y$ is obtained starting with the initial feasible solution $y:=0$ and incrementing some set of dual variables $\{y_{C}:v\in\mathcal{C}_{t}\}$ uniformly and maintaining feasibility of $y$ for iterations $t=1,2,..,l$ for some $l\in\mathbb{N}$.
  2. 2.
    For each iteration $t\in\{1,2,3,..,l\}$, the set $\{y_{C}:C\in\mathcal{C}_{t}\}$ of incremented dual variables satisfies $\sum_{C\in\mathcal{C}_{t}}|S\cap C|\leq\beta|\mathcal{C}_{t}|$.
  3. 3.
    $\forall v\in S,\ \ \sum_{C\in\mathcal{C}\ v\in C}y_{C}=c_{v}$.
Then $S$ has cost at most $\beta\sum_{C\in\mathcal{C}}y_{C}$, that is at most $\beta$ times the LP value.
Using Theorem 9, it suffices to prove that during any iteration $t$, the face minimal cycles $\mathcal{C}_{t}$ of $G^{S_{t}}$, where $S_{t}$ is our current hitting set satisfies
$\sum_{C\in\mathcal{C}_{t}}|\bar{S}\cap C|\leq O(g)|\mathcal{C}_{t}|.$ (2)
The following definition of crossing cycles was elementary to the approach by Goemans and Williamson [12].
Definition 2.
Fix an embedding of a planar graph. Two cycles $C_{1},C_{2}$ cross if $C_{i}$ contains an edge intersecting the interior of the region bounded by $C_{3-i}$, for $i=1,2$. That is, the plane curve corresponding to the embedding of the edge in the plane intersects the interior of the region of the plane bounded by $C_{3-i}$. A set of cycles $\mathcal{C}$ is laminar if no two elements of $\mathcal{C}$ cross.
Denote $\mathcal{C}^{\prime}$ the set of facial cycles of $\mathcal{C}$. For a node $v\in\bar{S}$, call a cycle $C\in\mathcal{C}^{\prime}$ with $C\cap\bar{S}$ a witness for $v$. Since we applied reverse deletion to $\bar{S}$ at the end of the algorithm, each node of $\bar{S}$ has a witness in $\mathcal{C}^{\prime}$ which is a cycle of $G^{S}$.
The following result about the structure of witness cycles was vital to the 3 and 2.4 approximations for DFVS in planar graphs by [12] and [3]. We observe that the proof in [12] which involves iteratively applying an “uncrossing” procedure to two witness cycles that cross yields the same result for facial cycles of graphs on surfaces.
Lemma 2.
[12] There exists a laminar family $\mathcal{A}\subset\mathcal{C}^{\prime}$ of witness cycles in $G^{S_{\bar{t}}}$ for $\bar{S}$.
The laminar family $\mathcal{A}$ can be represented by a forest where $A_{1}$ is an ancestor of $A_{2}$ if the inside region of $A_{1}$ contains the inside region of $A_{2}$. Add a root node $r$ to this forest, make it the parent of every maximal node of the forest and call the resulting tree $T$.
We assign each cycle $C$ of $\mathcal{C}_{t}$ to the smallest node of $T$ containing $C$. Call the set of cycles assigned to $w\in T$, $\mathcal{C}_{w}$. We assign the nodes that $w$ and the children of $w$ are witnesses of to $w$ and call this set $\bar{S_{w}}$.
For the purpose of bounding $\sum_{C\in\mathcal{C}_{t}}|\bar{S}\cap C|$, we define the following bipartite graph.
Definition 3.
[12] The debit graph for $\mathcal{C}_{t}$ and $S$ is the bipartite graph $\mathcal{D}_{G}=(\mathcal{R}\cup S,E)$ with edges $E_{\mathcal{C}_{t}}=\{(C,s)\in\mathcal{C}_{t}\times S\mid s\in C\}$.
Since each $C\in\mathcal{C}_{t}$ is incident to the vertices of $\bar{S}$ on $C$, $|\bar{S}\cap C|$ is the degree of $C$ in $\mathcal{D}_{G}$. Summing this equality over each $C\in\mathcal{C}_{t}$ yields $\sum_{C\in\mathcal{C}_{t}}|\bar{S}\cap C|=E(\mathcal{D}_{G})$. By placing the node of the debit graph corresponding to $C$ inside the inside region of $C$ we can see that the debit graph is also embedded on $Q$.
Proposition 10.
[Corollary of Euler’s formula for graphs of genus $g$] A (simple) bipartite graph $\bar{G}$ with at least three vertices embedded on a surface of genus $g$ satisfies
$E(\bar{G})\leq(2+g)|V(\bar{G})|-4$
if $G$ has two vertices then
$E(\bar{G})\leq(2+g)|V(\bar{G})|-3$
Proof.
Euler’s formula (for instance see [16]) for graphs embedded on a surface of genus $g$ yields $2-2g=|V(\bar{G})|-|E(\bar{G})|+|F(\bar{G})|$. Following the same method as the proof of Euler’s formula for bipartite planar graphs with at least 3 vertices, (for instance see Corollary 4.2.10 of [6]) we observe that each face of $\bar{G}$ having at least 4 edges means $|F(\bar{G})|\leq\frac{1}{2}||$. Thus for $|V(\bar{G})|\geq 3$, $|E(\bar{G})|\leq 2|V(\bar{G})|-4+4g\leq(2+g)|V(\bar{G})|-4$. If $|V(\bar{G})|\leq 2$ then $|E(\bar{G})|\leq 1\leq(2+g)|V(\bar{G})|-3$.
For a node $w$ of $T$ that is not a leaf or the root, the subgraph of $\mathcal{D}_{G}$ induced by $\mathcal{C}_{w}\cup\bar{S}_{w}$ is embedded on $Q$ and further $|\mathcal{C}_{w}\cup\bar{S}_{w}|\geq 3$, thus by 10,
$|E(\mathcal{D}_{G}(\mathcal{C}_{w}\cup\bar{S}_{w}))|\leq(2+g)|\mathcal{C}_{w}|+(2+g)|\bar{S}_{w}|-4=(2+g)|\mathcal{C}_{w}|+(2+g)(\operatorname{deg}_{T}(w)-1)-4.$ (3)
For a leaf $v$ of $T$
$|E(\mathcal{D}_{G}(\mathcal{C}_{v}\cup\bar{S}_{v}))|\leq(2+g)|\mathcal{C}_{v}|+2|\bar{S}_{v}|-3=(2+g)|\mathcal{C}_{v}|+2(\operatorname{deg}_{T}(v)-1)-3.$ (4)
For the root $r$ of $T$
$|E(\mathcal{D}_{G}(\mathcal{C}_{r}\cup\bar{S}_{r}))|\leq(2+g)|\mathcal{C}_{r}|+2|\bar{S}_{r}|=(2+g)|\mathcal{C}_{r}|+2(\operatorname{deg}_{T}(r)-1).$ (5)
Summing these up we get
$\begin{array}[]{ll}|E(\mathcal{D}_{G})|&=\sum_{v\in T}|E(\mathcal{D}_{G}(\mathcal{C}_{v}\cup\bar{S}_{v}))|\\ &\leq(2+g)|\mathcal{C}|+\sum_{v\in T}(2+g)\operatorname{deg}_{T}(v)-4|T|+l+4\\ &\leq(2+g)|\mathcal{C}|+2((2+g)|T|-2)-4|T|+l+4\\ &\leq(2+g)|\mathcal{C}|+2g|T|+l\\ &\leq(3+3g)|\mathcal{C}|\par\end{array}$
where $l$ is the number of (non-root) leaves of $T$. Thus, $\sum_{C\in\mathcal{C}_{t}}|\bar{S}\cap C|=|E(\mathcal{D}_{G})|\leq(3+3g)|\mathcal{C}|$.
This shows that $\bar{S}$ has cost $O(g)OPT_{LP}$ and hence our algorithm returns a solution of cost $O(g)OPT_{LP}$. ∎
Appendix A Statement and proofs of topological results we use
First let us prove 7.
Proof.
We use the following theorem.
Theorem 11.
([24] pg 308) In any minimal genus embedding of a graph, each face is homeomorphic to an open disk.
[…]
Case 1: At least one of $G_{L}$ or $G_{R}$ is empty.
- Suppose, without loss of generality, that $G_{L}$ is empty. Consider the face of $G$ that contains $C$ and intersects the left of $C$. By Youngs’s theorem, this face must be homeomorphic to a disc. But, $C$ is not contractible (else it would separate the surface $Q$ into two components). Hence, a small leftward shift of $C$ which will lie in the face $f$ will also not be contractible. This contradicts that $f$ is homeomorphic to a disc, since a disc is simply connected.
Case 2: Both $G_{L}$ and $G_{R}$ are nonempty.
- We first show that if a face contains vertices of $G_{L}$ and of $C$ (resp., of $G_{R}$ and of $C$), then it cannot contain vertices of $G_{R}$ (resp, of $G_{L}$). Suppose, for the sake of contradiction, that there is such a face $f$ of degree $d$. Let $\partial f=v_{0}\dotsm v_{d-1}$ be the boundary cycle of $f$, where $i\in\mathbb{Z}/d\mathbb{Z}$. Without loss of generality, assume that $v_{0}\in\hat{L}$ and for some $q$ $v_{1},v_{2},..,v_{q}\in C$, and $v_{q+1}\in\hat{R}$. There is are points $p_{l}$ on the edge $v_{0}v_{1}$ in the interior of $L$ and $p_{R}$ on the edge $v_{q}v_{q+r}$ in the interior of $R$. Let $h:[0,1]\rightarrow Q$ be a non-self-intersecting curve in $f$ from $p_{R}$ to $p_{L}$. Let $r_{L},r_{R}>0$ be such that $B_{Q}(p_{l},r_{L})\subset L,\ \ B_{Q}(p_{R},r_{R})\subset R$. Let $h^{L},h^{R}$ be non-self-intersecting curves in $B_{Q}(p_{l},r_{L})$ and $B_{Q}(p_{R},r_{R})$ respectively not intersecting $h$. Then $h\cup h^{L}\cup h^{R}$ satisfies the conditions of 4. Thus $h\cup h^{L}\cup h^{R}\cup g(p_{l}v_{0},v_{1},...,v_{q+1}p_{R})$ does not bound a region of the closure of $f$ homeomorphic to a disk which cannot happen because $f$ is homeomorphic to an open disk.
- Next, we show that if there is a face of $G$ that contains vertices of $G_{L}$ and of $G_{R}$ (but not of $C$), then we again arrive at a contradiction: the boundary of such a face $f$ will necessarily contain more than one connected component, since there can be no edge between vertices of $G_{L}$ and of $G_{R}$. But, this is not possible since $f$ is homeomorphic to a disc by Youngs’s theorem.
- Now, consider the subsets $Q_{L}$ and $Q_{R}$ of $Q$ obtained by taking the union of all the vertices, edges and faces induced by $G_{L}\cup C$ and $G_{R}\cup C$, respectively. By assumption 3, every component of $G-C$ is in $G_{L}$ or $G_{R}$, so every vertex and edge of $G$ belongs to $Q_{L}$ or $Q_{R}$. By the subcases eliminated above under Case 2, every face of $G$ also belongs to $Q_{L}$ or $Q_{R}$ (but not both). Then, $Q=Q_{L}\cup Q_{R}=(Q_{L}-C)\sqcup(Q_{R}-C)\sqcup C$. This means that $C$ separates $Q$ into two components, which contradicts assumption 1. ∎
It is well known (see for instance [1] page 15) that smooth surfaces $Q$ have the property that for each $v\in Q$ there is an open ball $B_{Q}(v,r_{0})$ of some small radius $r_{0}>0$ in $Q$ and a diffeomorphism $\psi$ from $B_{Q}(v,r_{0})$ to the open disk $B_{\mathbb{R}^{2}}(0,r_{0})$ of radius $r_{0}$ about the origin in the two-dimensional plane such that $\psi(v)=(0,0)$ and $\psi$ preserves distances from $v$, that is $\operatorname{dist}_{Q}(v,x)=\|\psi(v)-\psi(x)\|$, where $\operatorname{dist}_{Q}(v,x)$ is the geodesic distance from $v$ to $x$ in $Q$. For $p\in Q$, $r>0$, denote by $B(p,r)$ the open ball of radius $r$ about $p$. We now formally state and prove what 3 and 4 informally say.
Proposition 12.
Given a closed continuous non-self-intersecting curve $C^{\prime}$ embedded on an orientable surface $Q$. There exist some radius $r>0$ and disjoint subsets $L,R$ “on each side” of $C^{\prime}$ such that the set $\{B(v,r):\ v\in C^{\prime}\}$ (where $B(v,r)$ is the open ball around $v$ of radius $r$ in $Q$) is contained in the union $L\cup R\cup C^{\prime}$, and for each $v\in C^{\prime}$, $r^{\prime}\leq r$, $L\cap B(v,r^{\prime})$ and $R\cap B(v,r^{\prime})$ are the two connected components of $B(v,r^{\prime})\backslash C^{\prime}$. There is a diffeomorphism $\phi$ from $L\cup C^{\prime}\cup R$ to a connected open neighbourhood of $C^{\prime}\times\{0\}$ in $C^{\prime}\times\mathbb{R}$ and small $q>0$ with $C^{\prime}\times(-q,0)\subset\phi(L)\subset C^{\prime}\times(-\infty,0)$, $C^{\prime}\times(0,q)\subset\phi(R)\subset C^{\prime}\times(0,\infty)$ and $\phi(C^{\prime})=C^{\prime}\times\{0\}$. Further for any (piecewise smooth) curve $f:[0,1]\rightarrow Q$ such that $f(x)\notin C^{\prime}$ for any $x\in[0,1)$, $f(1)\in C^{\prime}$ satisfies that for some $\beta\in(0,1)$, either $f((\beta,1))\in L$, that is the curve “reaches $C$ from the left” $L$ or $f((\beta,1))\in R$, that is the curve “reaches $C^{\prime}$ from the right” $R$.
Proposition 13.
For a finite set of curves $f_{1},f_{2},f_{3},..f_{l^{\prime}},h_{1},h_{2},..,h_{t^{\prime}}:[0,1]\rightarrow Q$ such that for each $i$, $f_{i}(x)\notin C^{\prime}$ for any $x\in[0,1)$ and $h_{i}(x)\notin C^{\prime}$ for any $x\in[0,1]$ we may choose $L,R,r$ above so that each curve $f_{i}([0,1))$ is disjoint from at least one of $L,R$ and each curve $h_{i}([0,1])$ is disjoint from both $L,R$. We refer to $L$ and $R$ as the left and right of $C^{\prime}$ respectively.
Further, there are curves $f_{L}:[0,1]\rightarrow L$, $f_{R}:[0,1]\rightarrow R$ which are homotopic to $C^{\prime}$. Informally speaking, these are obtained by “slightly shifting” $f$ “left” and “right” respectively.
Proposition 14.
Lastly let $h:[0,1]\rightarrow Q$ be any curve that reaches $C^{\prime}$ from the right at a point $c_{2}=h(1)$ on $C^{\prime}$, leaves $C^{\prime}$ from the left at $c_{1}=h(0)$, that is the curve $\bar{\psi}(t)=h(1-t)$ reaches $C^{\prime}$ at $c_{1}$ from the left and $h$ is otherwise disjoint from $C^{\prime}$. Assume $c_{1}\neq c_{2}$ and let $C^{\prime}_{c_{1},c_{2}}$ be a subcurve of $C^{\prime}$ with endpoints $c_{1}$ and $c_{2}$. Then there is a curve $\hat{h}:[0,2]\rightarrow Q$ that reaches $C^{\prime}$ from the right at $c_{1}=\hat{h}(1)$ and leaves $C^{\prime}$ at a point $c_{1}=\hat{h}(0)$ $-^\hat{h}$ is otherwise disjoint from $C^{\prime}$, and a there is a homeomorphism of $Q$ that maps $-^\hat{h}$ to the concatenation of $h$ and $C^{\prime}_{c_{1},c_{2}}$.
Proposition 15.
Let $Q$ be an orientable surface and $\phi:[0,1]\rightarrow Q$ a closed curve not dividing $Q$ into 2 regions with disjoint subsets $L,R$ “on each side” of $\phi$ as in 12. Let $c_{1},c_{2}\in[0,1)$, with $c_{2}\geq c_{1}$. Suppose that $\phi_{1}:[0,1]\rightarrow Q$ is a curve with $\phi_{1}(0)=\phi(c_{1})\ \ \phi_{1}(1)=\phi(c_{2})$, $\phi_{1}((0,1))$ is disjoint from $\phi([0,1])$ and the curve $\phi_{1}([0,0.5])$ approaches $\phi([0,1])$ from the left $L$ and $\phi_{1}([0.5,1])$ approaches $\phi([0,1])$ from the right $R$.
Then the curve $\phi_{2}:[0,1]\rightarrow Q$, $\phi_{2}(x)=\phi(c_{2}-x)$ if $x\leq c_{2}-c_{1}$ and $\phi_{2}(x)=\phi_{1}(\frac{1}{c_{2}-c_{1}}(x-c_{2}+c_{1}))$ for $x>c_{2}-c_{1}$, that is the curve obtained by joining the portion of $\phi$ from $c_{1}$ to $c_{2}$ to $\phi_{1}([0,1])$ does not divide $Q$ into 2 regions.
Proof.
We prove 12, 13, and 14. We use the following corollary of the tubular neighbourhood theorem (see for instance [4]).
Proposition 16.
[corollary of tubular neighbourhood theorem [4]] Given a curve $C^{\prime}$ embedded in a surface $Q$ there is an open neighbourhood $U$ of $C$ and an open set $V$ in $C^{\prime}\times\mathbb{R}$ such that there is a diffeomorphism $\phi:U\rightarrow V$ with $\phi(C^{\prime})=C^{\prime}\times\{0\}$.
Let $w:[0,1]\rightarrow C^{\prime}$ with $w(0)=w(1)$, $w(0.5)=c_{2}$ be a parameterization of $C^{\prime}$. We define distance on $C^{\prime}\times\mathbb{R}$ by $\operatorname{dist}((a_{1},b_{1}),(a_{2},b_{2})):=(\operatorname{dist}_{Q}(a_{1},a_{2})^{2}+|b_{1}-b_{2}|^{2})^{\frac{1}{2}}$, where $\operatorname{dist}_{Q}(a_{1},a_{2})$ is the geodesic distance between $a_{1}$ and $a_{2}$ in $Q$. For each $v\in C^{\prime}$ let $q_{v}$ be the minimum of 1 and $\operatorname{sup}\{q^{\prime}:\ B(v,q^{\prime})\subset V\}$. $q_{v}$ is continuous in $v$ and $q_{v}>0\ \forall v\in C^{\prime}$. By compactness of $C^{\prime}$, $q:=\operatorname{min}_{v\in C^{\prime}}q_{v}$ exists and is positive. Now define $L^{\prime}=\phi^{-1}(C^{\prime}\times(-q,0))$, $R^{\prime}=\phi^{-1}(C^{\prime}\times(0,q)$.
Define $U^{\prime}=L^{\prime}\cup C^{\prime}\cup R^{\prime}$. Let $f:[0,1]\rightarrow Q$ be a curve with $f(x)\notin C^{\prime}$ for any $x\in[0,1)$. By continuity of $f$ there is some $\beta\in(0,1)$ for which $f((\beta,1])\subset U^{\prime}$. We claim $f((\beta,1))\subset L^{\prime}$ or $f((\beta,1])\subset R^{\prime}$. If $\phi(f((\beta,1)))$ contains a point in $C^{\prime}\times(-\infty,0)$ and a point in $\phi^{-1}(C^{\prime}\times(0,\infty)$ then by continuity $\phi(f((\beta,1)))$ contains a point in $C^{\prime}\times\{0\}$ and hence $f((\beta,1))$ contains a point in $C^{\prime}$ which is a contradiction. Thus either $\phi(f((\beta,1)))\subset C^{\prime}\times(-\infty,0)$ or $\phi(f((\beta,1)))\subset C^{\prime}\times(0,\infty)$. If $\phi(f((\beta,1)))\subset C^{\prime}\times(-\infty,0)$, then $f((\beta,1))\subset L^{\prime}$. If $\phi(f((\beta,1)))\subset C^{\prime}\times(0,\infty)$, then $f((\beta,1))\subset R^{\prime}$.
For each $f_{i}$ there exists $\beta_{i}$ for which $f((\beta_{i},1])\in L^{\prime}$ or $f((\beta_{i},1])\in R^{\prime}$. For each $x\in C^{\prime}$ define $r^{\prime}_{x}$ to be the supremum of all radius $r^{\prime\prime}_{x}$ for which the ball $B(x,r^{\prime\prime}_{x})$ of radius $r^{\prime\prime}_{x}$ is entirely contained in $U^{\prime}$ and for which $B(x,r^{\prime\prime}_{x})$ is disjoint from $f_{1}([0,\beta_{1})),f_{2}([0,\beta_{s})),f_{3}([0,\beta_{3})),..f_{l^{\prime}}([0,\beta_{l})),$ $h_{1}([0,1]),h_{2}([0,1]),..,h_{t^{\prime}}([0,1])$. If the supremum does not exist, set $r^{\prime}_{x}=\infty$. Define $r_{x}=\operatorname{min}\{1,r^{\prime}_{x}\}$. Again $r_{x}>0$ for all $x\in C^{\prime}$ and is continuous in $x$. Since $C^{\prime}$ is a compact set, $r:=\operatorname{min}_{x\in C^{\prime}}r_{x}$ exists and is positive. Note each curve $f_{i}([0,1))$ is disjoint from at least one of $L^{\prime}\cap\{B(x,r):\ x\in C\},R^{\prime}\cap\{B(x,r):\ x\in C\}$ and each curve $h_{i}([0,1])$ is disjoint from both $L^{\prime}\cap\{B(x,r):\ x\in C\},R^{\prime}\cap\{B(x,r):\ x\in C\}$.
Let us show that by making $r$ smaller if necessary $B(v,r)\backslash C^{\prime}$ contains two connected components.
Proposition 17.
Given a (piece-wise smooth non-self-intersecting) curve $f:[0,1]\rightarrow\mathbb{R}^{2}$ with $t_{0}\in(0,1)$ there exists $r>0$ for which $B(f(t_{0}),r)\backslash f([0,1])$ contains exactly two components.
Proof.
Let $[t_{0},t_{1}]$ be an interval in which $f$ is smooth.
Let $f(t)=f(t_{0})+(t-t_{0})\nabla f(t_{0})+g(t-t_{0})$. By smoothness of $f$, $\nabla g(t-t_{0})$ is bounded for $t\in[t_{0},t_{1}]$ and $g(t-t_{0})=o(t-t_{0})$. Differentiating $\|f(t)-f(t_{0})\|^{2}=\|(t-t_{0})\nabla f(t_{0})+g(t-t_{0})\|^{2}$ we obtain
$\frac{d}{dt}\|f(t)-f(t_{0})\|^{2}=2(\nabla f(t_{0})+\nabla g(t-t_{0}))^{t}((t-t_{0})\nabla f(t_{0})+g(t-t_{0}))$
$=2(\nabla f(t_{0})+o(1))^{t}((t-t_{0})\nabla f(t_{0})+o(t-t_{0}))$
$=2(\nabla f(t_{0})+o(1))^{t}(t-t_{0})\nabla f(t_{0})+o(t-t_{0})$
For $t$ close enough to $t_{0}$ the last line is positive. This implies that for some $t_{2}>t_{0}$, $\|f(t)-f(t_{0})\|^{2}$ is increasing on $[t_{0},t_{2}]$. Likewise for some $t_{3}<t_{0}$, $\|f(t)-f(t_{0})\|^{2}$ is decreasing on $[t_{3},t_{0}]$.
Let $r>0$ be such that $\|f(t_{0})-f(t)\|\geq 2r$ for all $t\in[0,1]\backslash\operatorname{frac}([t_{3},t_{2}])$, where $\operatorname{frac}(x)$ is the fractional part of $x$. Then $\|f(t_{0})-f(t_{2})\|,\|f(t_{0})-f(t_{3})\|\geq 2r$. Then since $\|f(t_{0})-f(t)\|$ is increasing on $[t_{0},t_{2}]$ there is exactly one $t_{4}\in[t_{0},t_{2}]$ with $\|f(t_{0})-f(t_{4})\|=r$. Likewise there is exactly one $t_{5}\in[t_{3},t_{0}]$ with $\|f(t_{0})-f(t_{5})\|=r$. So $f([t_{5},t_{4}])$ forms a simple curve in the closed ball $\bar{B}(f(t_{0}),r)$ with endpoints on the boundary and $f((t_{5},t_{4}))$ lying in the interior. It follows from the Jordan curve theorem that $B(f(t_{0}),r)\backslash f([0,1])=B(f(t_{0}),r)\backslash f((t_{5},t_{4}))$ contains exactly two connected components. ∎
Let $v\in C^{\prime}$. For small enough $r_{0}$, $B(v,r_{0})$ is diffeomorphic to the open disk $B(0,r_{0})$ in $\mathbb{R}^{2}$ via some diffeomorphism $\psi$ with $\psi(v)=(0,0)$. Let $w$ be a paramaterization of $C^{\prime}$ with $w(0.5)=v$. Let $t_{1}:=\operatorname{inf}\{t:\ w([t,0.5])\subset B(v,r_{0})\}$ and $t_{2}:=\operatorname{sup}\{t:\ w([0.5,t])\subset B(v,r_{0})\}$ that is $[t_{1},t_{2}]$ is a maximal interval for which $w([t_{1},t_{2}])\subset B(v,r_{0})$ by continuity $t_{1}<0.5<t_{2}$. Choose $r_{1}>0$ less than the distance from $v$ to $C^{\prime}\backslash w((t_{1},t_{2}))$ and $r_{1}<\operatorname{dist}_{Q}(v,w(t_{1})),\operatorname{dist}_{Q}(v,w(t_{2}))$. From the previous proposition there exists $r_{2}>0$ such that $B_{\mathbb{R}^{2}}(\psi(v),r_{2})\backslash\psi(w([t_{1},t_{2}]))$ contains exactly two connected components. By making $r$ smaller than $r_{1}$ and $r_{2}$ if necessary we get that for any $0<\hat{r}\leq r$, $B_{\mathbb{R}^{2}}(\psi(v),\hat{r})\backslash\psi(C^{\prime})$ contains exactly two connected components. Thus, $B(v,\hat{r})\backslash C^{\prime}$ contains exactly two connected components.
Define $L=L^{\prime}\cap\{B(x,r):\ x\in C\}$, $R=R^{\prime}\cap\{B(x,r):\ x\in C\}$. Each curve $f:[0,1]\rightarrow Q$ with $f(x)\notin C^{\prime}$ for any $x\in[0,1)$ satisfies $f((\beta,1))\subset L$ or $f((\beta,1])\subset R$. Each curve $f_{i}([0,1))$ is disjoint from at least one of $L,R$ and each curve $h_{i}([0,1])$ is disjoint from both $L,R$.
Since $\phi(v)=\{v\}\times\{0\}$, $\phi(B(v,\hat{r}))$ intersects both $\{v\}\times(-\infty,0)$ and $\{v\}\times(0,\infty)$. Recall $B(v,\hat{r})\subset L\cup R\cup C^{\prime}$, so $B(v,\hat{r})$ intersects both $L$ and $R$. Since there is no path from $L$ to $R$ in $L\cup R\cup C^{\prime}$ one component of $B(v,\hat{r})\backslash C^{\prime}$ is contained in $L$ and the other is contained in $R$.
For each $v\in C^{\prime}$ let $y_{v}$ be the supremum of $\{y^{\prime}_{v}\geq 0:\ \{v\}\times(-y^{\prime}_{v},y^{\prime}_{v})\subset\phi(B(v,r))\}$. Again $y_{v}$ is positive and continuous in $v$ so $y:=\operatorname{min}_{v\in C^{\prime}}y_{v}$ exists and is positive. Then $C^{\prime}\times(-y,0)\subset\phi(L)$ and $C^{\prime}\times(0,y)\subset\phi(R)$. Define the curves $f_{L},f_{R}$ to be parameterizations of $\phi^{-1}(C^{\prime}\times\{\frac{-y}{2}\})$ and $\phi^{-1}(C^{\prime}\times\{\frac{y}{2}\})$ respectively.
Lastly given a curve $h:[0,1]\rightarrow Q$ be any curve that reaches $C^{\prime}$ from the right at a point $c_{2}=h(1)$ on $C^{\prime}$, leaves $C^{\prime}$ from the left at $c_{1}=h(0)$ and $C^{\prime}_{c_{1},c_{2}}$ be a subcurve of $C^{\prime}$ with endpoints $c_{1}$ and $c_{2}$. Let $j:[0,1]\rightarrow C^{\prime}_{c_{1},c_{2}}$ be a parameterization of $C^{\prime}_{c_{1},c_{2}}$ and denote $\bar{j}:[0,2]\rightarrow Q$ by $\bar{j}(t)=h(t)$ for $t\in[0,1]$ and $\bar{j}(t)=j(t-1)$ for $t>1$. Informally speaking, we “slightly shift” all points in $L\cup C^{\prime}\cup R$ to the right while keeping $h([0,1])\cap L$ to the left of $C^{\prime}$. Let $\phi(x)=(\phi_{1}(x),\phi_{2}(x))$, $\phi^{-1}(x)=(\phi^{-1}_{1}(x),\phi^{-1}_{2}(x))$.
Let $\gamma:C^{\prime}\rightarrow(-y,0)$ be any continuous function such that $\gamma(c_{2})=0$ and for any $t\in[0,1]$ for which $\phi(h(t))\in C^{\prime}\times(-y,0)$, $(\phi(h(t))_{2})<\gamma(\phi(h(t))_{1})<0$. Informally $\gamma$ is a curve lying to the right of $\phi(h([0,1]))\cap C^{\prime}\times(-y,0)$ and to the left of $[0,1]\times\{0\}$. Such $\gamma$ exists for instance define $-\gamma(t)$ to be half of the minimum of the distance from the point $(t,0)$ to $\phi(h([0,1]))\cap C^{\prime}\times(-y,0)$ and $y$.
Define $\bar{\gamma}:C^{\prime}\times(-y,y)\rightarrow C^{\prime}\times(-y,y)$ as follows. For $(a,b)\in C^{\prime}\times(-y,y)$ if $b<\frac{\gamma(a)}{4}$ define $\bar{\gamma}(a,b)=(a,b)$. If $\frac{\gamma(a)}{4}\leq b<0$, define $\bar{\gamma}(a,b)=(a,\frac{\gamma(a)}{4}+2(b-\frac{\gamma(a)}{4}))$. If $0\leq b\leq\frac{-\gamma(a)}{2}$, define $\bar{\gamma}(a,b)=(a,\frac{-\gamma(a)}{4}+\frac{b}{2})$. If $b\geq\frac{-\gamma(a)}{2}$ $\bar{\gamma}(a,b)=(a,b)$. Informally, $\bar{\gamma}$ shifts $C^{\prime}\times(-y,y)$ to the right while keeping $\phi(h([0,1]))\cap C^{\prime}\times(-y,0)$ left of $C^{\prime}\times\{0\}$.
Define $\hat{\gamma}:Q\rightarrow Q$ by $\gamma$ by $\hat{\gamma}(v)=v$ if $v\notin\phi^{-1}(C^{\prime}\times(-y,y))$ and $\hat{\gamma}(v)=\phi^{-1}\bar{\gamma}(\phi(v))$ if $v\in\phi^{-1}(C^{\prime}\times(-y,y))$. Note $\hat{\gamma}$ is a homeomorphism from $\phi^{-1}(C^{\prime}\times(-y,y))$ to $\phi^{-1}(C^{\prime}\times(-y,y))$ and from $Q\backslash\phi^{-1}(C^{\prime}\times(-y,y))$ to $Q\backslash\phi^{-1}(C^{\prime}\times(-y,y))$. Further, $\hat{\gamma}$ agrees on the boundary of $\phi^{-1}(C^{\prime}\times(-y,y))$ and $Q\backslash\phi^{-1}(C^{\prime}\times(-y,y))$, that is for sequence $\{a_{i}\}_{i=1}^{\infty}$ converging to a boundary point $a$ of $\phi^{-1}(C^{\prime}\times(-y,y))$ (resp $Q\backslash\phi^{-1}(C^{\prime}\times(-y,y))$) the sequence $\{\hat{\gamma}(a_{i})\}_{i=1}^{\infty}$ converges to $\hat{\gamma}(a)$. Thus, $\hat{\gamma}$ is a homeomorphism on $Q$, in fact, it turns out to be a continuous deformation.
Define $\hat{h}:[0,2]\rightarrow Q$ by $\hat{h}(t)=\hat{\gamma}(\bar{j}(t))$. Then $\hat{\gamma}$ is a homeomorphism mapping $\hat{h}([0,2])$ to $\bar{j}([0,2])$.
The actual statement of the tubular neighbourhood theorem involves first defining the normal fibre $N_{x}$ as the quotient $T_{x}Q/T_{x}C$ where $T_{x}Q$ and $T_{x}C$ are the tangent plane and tangent to the curve $C$ at $x$ and the normal bundle $NX$ as $\{(x,v):\ x\in C\ v\in N_{x}\}$.
Theorem 18 (tubular neighbourhood theorem).
There are open sets $U$ in $Q$ containing $C$ and $V$ in $NX$ such that there is a diffeomorphism $\gamma:U\rightarrow V$.
Let us quickly show how the version of the tubular neighbourhood theorem in 16 follows from the tubular neighbourhood theorem.
By orientability each point $x\in Q$ has a normal vector $n(x)$ and $n$ is continuous. We may parameterize $C^{\prime}$ with a function $\psi:[0,\beta]\rightarrow C^{\prime}$ with derivative $\psi^{\prime}(x)=1$ for some $\beta$. Define $v(x)$ to be of unit norm and positively orthogonal to $n(x),\phi^{\prime}(x)$ that is $n(x)^{t}v(x)=(\phi^{\prime}(x))^{t}v(x)=0$ and $\begin{bmatrix}n(x)&\psi^{\prime}(x)&v(x)\end{bmatrix}$ has determinant 1. By the inverse function theorem, $v(x)$ is continuous. $n(x),\psi^{\prime}(x),v(x)$ is a basis for $\mathbb{R}^{3}$ known as the Darboux frame. Since $v(x)$ is orthogonal to $n(x)$ it lies in the tangent plant $T_{x}Q$ since $v(x)$ is orthogonal to $\phi^{\prime}(x)$, $v(x),\phi^{\prime}(x)$ is a basis for $T_{x}Q$. Thus, $N_{x}=T_{x}Q/T_{x}C$ is diffeomorphic to $\{av(x):\ a\in\mathbb{R}\}$ which is diffeomorphic to $\mathbb{R}$. Thus, $NX$ is diffeomorphic to $C^{\prime}\times\mathbb{R}$.
To prove 15, we first prove the following special case.
Proposition 19.
Let $Q$ be an orientable surface and $\phi:[0,1]\rightarrow Q$ a closed curve not dividing $Q$ into 2 regions with disjoint subsets $L,R$ “on each side” of $\phi$ as in 12. Let $c\in[0,1)$, with $c_{2}\geq c_{1}$. Suppose that $\phi_{1}:[0,1]\rightarrow Q$ is a curve with $\phi_{1}(0)=\phi(c)\ \ \phi_{1}(1)=\phi(c)$, $\phi((0,1))$ is disjoint from $\phi([0,1])$ and the curve $\phi_{1}([0,0.5])$ approaches $\phi([0,1])$ from the left $L$ and $\phi_{1}([0.5,1])$ approaches $\phi([0,1])$ from the right $R$.
Then $\phi_{1}$ does not divide $Q$ into 2 regions.
Proof.
Suppose for a contradiction that $\phi_{1}$ divides $Q$ into 2 regions. It’s a well-known result that one of the regions $Q_{1}$ must be homeomorphic to an open disk. Let $Q_{2}$ be the other region.
There exists a small radius $r_{1}$ for which the ball $B(\phi(c),r)$ of radius $r_{1}$ about $\phi(c)$ such that $B(\phi(c),r_{1})$ is homeomorphic to an open disk. Let $r$ be as in 12 and define $r_{2}=\operatorname{min}\{r,r_{1}\}$.
Definition 4.
Given 2 curves $f_{1},f_{2}:[0,1]\rightarrow Q$ on a surface $Q$, with $f_{1}(0.5)=f_{2}(0.5)$ and $f_{1}(x)\neq f_{2}(y)\ \ \forall x,y\in([0,1]\backslash\{0.5\})$, that is they intersect only at $f_{1}(0,5)$, we say that $f_{1}$ crosses $f_{2}$ at $f_{1}(0.5)$ if there exists $r_{0}>0$ such that for all $r\leq r_{0}$ such that $f_{2}$ intersects both regions of $B(f_{1}(0.5),r)\backslash f_{1}([0,1])$, where $B(p,r)$ is the open ball around $p$ of radius $r$.
Lemma 3.
For two curves $f_{1},f_{2}$ on a surface $Q$ and $p$ a point on both curves, $f_{1}$ crosses $f_{2}$ at $p$ if and only if $f_{2}$ crosses $f_{1}$ at $p$.
Proof.
Let $f_{1}(b_{1})=p=f_{2}(b_{2})$. Let $L_{1}$, $R_{1}$ and $r_{1}$ (resp $L_{2}$, $R_{2}$ and $r_{2}$) be the left, right and radius respectively for $f_{1}$ (resp. $f_{2}$ ) as guaranteed by 12. Define $r=\operatorname{min}\{r_{1},r_{2}\}$.
Claim 5.
$f_{1}$ crosses $f_{2}$ at $p$ if and only if for all $r\geq r_{0}>0$ none of $L_{1}\cap B(p,r_{0})$, $R_{1}\cap B(p,r_{0})$, $L_{2}\cap B(p,r_{0})$, $R_{2}\cap B(p,r_{0})$ is contained in another.
Proof.
Suppose that $f_{1}$ crosses $f_{2}$. Let $r\geq r_{0}>0$. Let $t_{L},t_{R}\in\mathbb{R}$ be such that $f_{2}(t_{L})\in L_{1}\cap B(p,r_{0}),\ f_{2}(t_{R})\in R_{1}\cap B(p,r_{0})$. Since $R_{1}\cap B(p,r_{0}),L_{1}\cap B(p,r_{0})$ are open, there exists $r_{L},r_{R}$ be such that $B(f_{2}(t_{L}),r_{L})\subset L_{1}\cap B(p,r_{0})$ and $B(f_{2}(t_{R}),r_{R})\subset R_{1}\cap B(p,r_{0})$. Since $B(f_{2}(t_{L}),r_{L})\cap L_{2},B(f_{2}(t_{L}),r_{L})\cap R_{2},B(f_{2}(t_{R}),r_{R})\cap L_{2},B(f_{2}(t_{R}),r_{R})\cap R_{2}\neq\emptyset$, none of $L_{1}\cap B(p,r_{0})$, $R_{1}\cap B(p,r_{0})$, $L_{2}\cap B(p,r_{0})$, $R_{2}\cap B(p,r_{0})$ is contained in another.
Conversely, suppose that for any $0<r_{0}\leq r$ none of $L_{1}\cap B(p,r_{0})$, $R_{1}\cap B(p,r_{0})$, $L_{2}\cap B(p,r_{0})$, $R_{2}\cap B(p,r_{0})$ is contained in another. Then for any $0<r_{0}\leq r$, if $f_{2}$ does not intersect $L_{1}\cap B(p,r_{0})$, then $L_{1}\cap B(p,r_{0})$ is connected in $B(p,r_{0})\backslash f_{2}$, that is $L_{1}\cap B(p,r_{0})$ is contained in one of the two components $L_{2}\cap B(p,r_{0})$, $R_{2}\cap B(p,r_{0})$ of $B(p,r_{0})\backslash f_{2}$. Hence $f_{2}$ intersects $L_{1}\cap B(p,r_{0})$, likewise $f_{2}$ intersects $R_{1}\cap B(p,r_{0})$. Hence $f_{2}$ crosses $f_{1}$.
From the previous claim it’s clear that crossing is a symmetric relation. ∎
Define $\phi_{2}(x)=\phi(\operatorname{frac}(x-0.5+c))$, where $\operatorname{frac}(x)=x-\lfloor x\rfloor$ is the fractional value of $x$ and $\phi_{3}(x)=\phi_{1}(\operatorname{frac}(x-0.5))$, that is, $\phi_{2},\phi_{3}$ are reparameterized versions of $\phi$ and $\phi_{1}$. For some $10<\beta_{1}<\beta_{2}<$ 1 $\phi_{1}(x)\in L$ for $x\in(0,\beta_{1})$ and $\phi_{1}(x)\in R$ for $x\in(\beta_{2},1)$.
Define $\beta_{1}^{\prime}=\operatorname{min}\{\beta_{1},0.5\}+0.5$ $\beta_{2}^{\prime}=\operatorname{max}\{\beta_{2},0.5\}$. Then $\phi_{3}((0.5,\beta_{1}^{\prime}))\subset L$ and $\phi_{3}((\beta_{2}^{\prime},0.5))\subset R$. Since $L\cup R$ covers $\bar{B}(\phi(c),r)\backslash\phi([0,1])$, this implies that $\phi_{3}$ crosses $\phi_{2}$.
By 3 $\phi_{2}$ crosses $\phi_{3}$. Let $L_{\phi_{3}},R_{\phi_{3}}$ be the left and right of $\phi_{3}$ as in 12. Since $L_{\phi_{3}},R_{\phi_{3}}$ are connected, $L_{\phi_{3}},R_{\phi_{3}}$ belong to different regions of $Q\backslash\phi_{3}$. Since $\phi_{2}$ crosses $\phi_{3}$, there exists $t_{0}$ for which $\phi(t_{0})\in Q_{1}$ and $t_{1}$ for which $\phi(t_{1})\in Q_{2}$. Let $t_{2}=\operatorname{inf}\{a\in[0,1]:\ \exists b\in(a,1]\ \text{s.t.}\ \phi((a,b))\subset Q_{1}\}$, $t_{3}=\operatorname{sup}\{b\in[t_{2},1]:\ \text{s.t.}\phi((t_{2},b))\subset Q_{1}\}$, that is $[t_{2},t_{3}]$ is a maximal interval for which $\phi([t_{2},t_{3}])\subset Q_{1}$. It follows $\phi(t_{2}),\phi(t_{3})$ lie on the boundary of $Q_{1}$, that is on $\phi_{1}$. This implies $t_{2},t_{3}\in\{0,1\}$. Since $t_{2}<t_{3}$ $t_{2}=0$ and $t_{3}=1$. This implies that $\phi([0,1]))$ lies in $Q_{1}\cup\{\phi(0)\}$ contradicting that there exists $t_{1}$ for which $\phi(t_{1})\in Q_{2}$.
Proof.
(of 15) Let $f:[0,1]\rightarrow C^{\prime}$ be a paramaterization of $C^{\prime}$ and let $c=f^{-1}(c_{1})$. Note that by 14 the curve $\phi_{2}$ is homeomorphic to a curve $\phi_{3}$ that enters $\phi$ from the right and leaves $\phi$ from the left at the same point $f(c)$. By 19, $\phi_{3}$ does not divide $Q$ into 2 regions. Thus, $\phi_{2}$ does not divide $Q$ into 2 regions.